Identify the coordinates of the vertex, directrix, and the direction of opening of the parabola with equation (y-4)^2=x+2 Answer by lwsshak3(11628) 
 Identify the coordinates of the vertex, directrix, and the direction of opening of the parabola with equation (y-4)^2=x+2
**
This is an equation of a parabola that opens rightwards.
Its standard form: (y-k)2=4p(x-h)
For given equation:(y-4)^2=x+2
vertex: (-2,4)
Axis of symmetry:y=4
4p=1
p=1/4
directrix:x=-9/4 (p distance to the left of vertex on the axis of symmetry)